# Partial Knowledge Distillation for Alleviating the Inherent Inter-Class Discrepancy in Federated Learning

### v1

@@ Abstract
Substantial efforts have been devoted to alleviating the impact of the long-tailed class distribution in federated learning. In this work, we observe an interesting phenomenon that weak classes consistently exist even for class-balanced learning. These weak classes, different from the minority classes in the previous works, are inherent to data and remain fairly consistent for various network structures and learning paradigms. The inherent inter-class accuracy discrepancy can reach over 36.9% for federated learning on the FashionMNIST and CIFAR-10 datasets, even when the class distribution is balanced both globally and locally. In this study, we empirically analyze the potential reason for this phenomenon. Furthermore, a partial knowledge distillation method is proposed to improve the model’s classification accuracy for weak classes. In this approach, knowledge transfer is initiated upon the occurrence of specific misclassifications within certain weak classes. Experimental results show that the accuracy of weak classes can be improved by 10.7%, reducing the inherent inter-class discrepancy effectively.
Index Terms:
@@
Deep neural network, federated learning, class imbalance, weak class
I Introduction
@@
Federated learning (FL) is a decentralized machine learning paradigm where multiple clients collaborate to train a shared global model without directly sharing their local data [1]. This approach preserves data privacy by keeping the data on local devices. However, one big challenge to FL is the varying performance across different classes due to the presence of unbalanced sample sizes [2]. When the number of samples for different classes varies substantially, the global model often exhibits bias towards the majority classes, known as ‘head classes’. The minority classes, referred as ‘tail classes’, become underrepresented [3], thus leading to lower classification accuracy as depicted in Fig. 1(a). However, in this work, we show that ‘dominant’ and ‘weak’ classes consistently exist even when sample sizes are equalized and samples from each class are uniformly distributed across the clients (i.e. with a balanced class distribution both globally and locally). Unlike the minority classes discussed in prior studies, these weak classes are inherent to the data itself, not dictated by sample size. The resulting inherent inter-class accuracy discrepancy (ICD) is illustrated in Fig. 1(b).
@@
Fig. 2 presents the class-wise accuracy across several widely used benchmarks in the domain of FL. Each experiment is repeated five times, and the average result is presented. The ICD phenomenon identified in this study exhibits the following characteristics. 1) This phenomenon is universally present across different learning paradigms and especially pronounced in FL compared to centralized learning. 2) Models trained on the same dataset tend to exhibit a very similar ICD, indicating that this phenomenon is irrelevant to the network structure. 3) This phenomenon persists no matter whether the local class distributions in FL is balanced or not. Notably, as local imbalance intensifies, particularly when samples from weak classes are concentrated within a limited number of clients, the inherent ICD becomes significantly exacerbated. Details will be elaborated upon in the subsequent sections.
@@
Figure 1: (a) The Conventional Long-tailed Problem; (b) Inherent Inter-class Discrepancies Observed under Balanced Class Distributions (both Globally and Locally).
@@
Figure 2: (a) The Class-wise Accuracy (with Mean Subtraction) Based on Different Learning Paradigms and Network Structures ($\Delta$: The Maximum Discrepancy in One Training); (b) Visualization of the Class Distribution Among 10 Clients.
@@
Although promising progress has been made in alleviating the inter-class discrepancy brought by class imbalance [3], these techniques cannot be directly applied to the inherent ICD problem. Specifically, current approaches usually focus on rebalancing class distributions by employing techniques such as over-sampling [4] or data augmentation [5] for the minority classes. However, considering that the inherent ICD is not associated with the quantity of samples, these methods may not be sufficient to mitigate the problem effectively.
@@
Experiments indicate that a likely reason for this ICD phenomenon is the inherent similarity among the high-level feature representations of certain classes. This similarity poses a challenge for the classifier in accurately distinguishing data samples across these classes. In this work, we propose a federated partial knowledge distillation (PKD) method to improve classification accuracy for weak classes. Our method is inspired by the observation that an expert trained specifically on a group of confounding weak classes show greater proficiency in differentiating among them. This specialized model focuses solely on extracting and distinguishing the subtle feature differences among these classes and thus yields better accuracy than the model trained jointly on all classes. The term ‘partial’ in PKD is shown in several aspects: 1) An expert’s knowledge is confined to specific classes. It is tailored to identify a subset of data that shares analogous high-level features. 2) Knowledge distillation (KD) is triggered only when misclassifications between certain classes happens. Therefore, only a subset of the misclassified samples from these weak classes is considered in computing the KD loss during the training phase. 3) While there might be multiple experts, only up to one expert can be involved in the KD process for any given misclassified sample. Experimental results demonstrate that the proposed PKD method can effectively alleviate the inherent ICD issue by enhancing the classification accuracy of the worst performing classes. In summary, this work makes the following contributions:
@@
  • We highlight an interesting phenomenon that weak classes are always present, even in ideal scenarios where class distributions are balanced at both the global and local levels. This inherent variance in accuracy across different classes is especially pronounced in the context of FL, reaching up to 45.4% on the FashionMNIST dataset [6] and 36.9% on the CIFAR-10 dataset [7], respectively.
@@
  • Experimental results show that the proposed method can effectively improve the classification accuracy of the weak classes by over 10.7% and 6.6% on FashionMNIST and CIFAR-10, respectively.
II Inter-Class Accuracy Discrepancy
II-A Discrepancy Induced by the Class Imbalance Problem
@@
ICD due to class imbalance. A primary reason for the variance in class-wise accuracy in deep learning is the imbalanced class distribution [8, 9]. The unbalanced sample sizes lead to a dual-fold challenge: 1) Models tend to be biased towards the majority classes, leaving the minority classes under-represented. 2) The lack of data for minority classes makes it even harder to learn the characteristics of these classes [3]. In federated learning, the potential causes for ICD become more complex. The presence of class imbalance at either the global level or within the local dataset of individual clients may diminish the model’s quality. Besides, the mismatch in class distributions between the cloud and the clients may also degrade the performance of the global model [2]. Later we will show that the inherent ICD exits regardless of whether the local class distributions are balanced.
@@
Existing approaches. Existing solutions can be broadly classified into several categories. 1) Class re-balancing techniques: These methods focus on offsetting the negative impact of uneven class distribution in training samples. This can be achieved by either over-sampling the minority classes to ensure a more equitable presence in the training data [4, 10], or by re-weighting the training loss values to mitigate the influence of uneven positive gradients [11, 12]. 2) Data augmentation techniques: These methods seek to enlarge the sample size or enhance the sample quality of minority classes through data augmentation techniques [13, 14, 15]. Compared to mere re-balancing methods, this approach can potentially enrich the diversity within the minority classes, thereby reducing the risk of overfitting. 3) Other methods include meta-learning [16], ensemble learning [17, 18], transfer learning [17], and so on. Although significant progress has been achieved in reducing the ICD induced by class imbalance, such strategies may not be directly applicable to addressing the inherent ICD problem, as we will discuss below.
II-B The Inherent ICD in This Work
@@
Differences from the class imbalance problem. While the weak classes show similar poor accuracy to the minority classes in the class imbalance problem, their underlying causes and potential solutions are distinct. First, the inherent ICD is not a result of uneven sample sizes across different classes. Weak classes consistently exist even in the ideal case where both global and local data are class balanced. Consequently, previous methods that rely on re-sampling local data [14] or clients [19] based on their sample sizes are not feasible in this context. Second, in contrast to minority classes, weak classes are not characterized by data scarcity or a lack of diversity. Experimental results suggest that the conventional data augmentation techniques [20] fail to address the inherent ICD effectively, and in some cases, they might even intensify the inherent accuracy discrepancy.
Figure 3: FashionMNIST: (a) Raw Samples and the Similarity in High-Level Feature Representations between Each Pair of Classes; (b) Output Probabilities (Averaged over 1000 Samples per Class); (c) Prediction Results. For simplicity, each class is assigned a serial number. (0: T-shirt, 1: Trouser, 2: Pullover, 3: Dress, 4: Coat, 5: Sandal, 6: Shirt, 7: Sneaker, 8: Bag, 9: Ankle Boot.)
Figure 4: The Maximum and Minimum Accuracy During Training. (K: Number of Clients Participating in FL at each Round)
@@
The inherent ICD phenomenon. Fig. 2(a) shows the class-wise accuracy achieved on three datasets. In each scenario, the total number of samples for each class is kept identical, ensuring that global class balance is maintained. We adopt several commonly used convolutional neural networks and train the models following different learning paradigms. Two local data partition strategies are adopted in FL for each dataset, as illustrated in Fig. 2(b). Our results indicate that this inherent ICD is a common phenomenon across various network structures, learning paradigms, and local dataset partition strategies. Most of the dominant classes and the weak classes, e.g., {‘0’, ‘1’} and {‘8’, ‘9’} in MNIST [21], remain fairly consistent across various settings. The inter-class accuracy discrepancy is less marked for centralized learning and thus may go unnoticed. However, when it comes to FL, the discrepancy may increase significantly, e.g. reaching up to 45.4% for FashionMNIST. We also test the local class imbalance scenario using a pathological data partition strategy proposed in [1]. Specifically, each client is assigned samples from a single class to ensure a clear and straightforward comparison of the class-wise accuracy. As shown in Fig. 2(a), the same group of weak classes is observed, with the accuracy discrepancy further increasing to a significant 53.0% on the MNIST dataset. For more complex datasets like FashionMNIST and CIFAR-10, such discrepancy can reach up to over 73.9%.
@@
Fig. 4 shows the maximum and minimum class-wise accuracy at each round. While the overall accuracy rises smoothly over time, the difference between the maximum and minimum values continues to be evident. With a balanced class distribution globally and locally, the accuracy discrepancy on the FashionMNIST dataset is 28.9% when there are 10 clients selected at each round. This value increases significantly as the number of clients participating in FL grows to 100. In scenarios with local class imbalance, samples from weak classes are concentrated within a small subset of clients, thereby amplifying the accuracy discrepancy to 91.2%. The minimum class-wise accuracy is close to zero, indicating that the final model might completely fail in specific classes.
Potential reasons for inherent ICD. Fig. 3 (c) maps the predictions to the original classes to which the data belongs (the models are trained once). It shows that the negative samples of a weak class are often assigned certain wrong labels. For the FashionMNIST dataset with a balanced local class distribution, 62% (=$\frac{85}{137}$) of the misclassified ‘T-shirt’ samples are identified as ‘shirts’ using the LeNet-5 model [21]. Conversely, around 42% (=$\frac{132}{316}$) of the misclassified ‘shirt’ samples are labeled as ‘T-shirts’. Another group of classes that can easily be confused with each other include {‘pullover’, ‘coat’, and ‘shirt’}. For instance, 23% of the misclassified ‘shirt’ samples are identified as ‘pullovers’, while 23% are mistaken for ‘coats’. That is to say, a total of 88% of the misclassified ‘shirt’ samples are incorrectly categorized as other weak classes. This phenomenon becomes particularly pronounced in scenarios with local class imbalance. Approximately 81% of the overall samples categorized under ‘pullover’ and ‘coat’ are misidentified as ‘shirt’ by the LeNet-5 model. A similar phenomenon is also observed with ConvNet, as shown in Fig. 3 (c).
Fig. 3(a) presents some misclassified samples from the two groups of weak classes, i.e. {‘T-shirt’, ‘shirt’} and {‘pullover’, ‘coat’, ‘shirt’}. It can be seen that these samples exhibit considerable similarities. Fig. 3(a) assesses the similarity in the high-level features between each pair of classes. The weak classes within the two aforementioned groups exhibit a lower Euclidean distance compared to the others, indicating a higher degree of similarity. As a result, it is challenging for the classifier to differentiate among these classes, leading to lower accuracy and reduced confidence in the model’s predictions. The probability distributions determined by the softmax function are depicted in Fig. 3(b). The figure displays the average probability distribution across all samples within each class.
We further investigate how the classification accuracy for weak classes changes during training. As depicted in Fig. 5, the variance between two consecutive rounds can reach up to 80%, even when the global accuracy has become relatively stable. Furthermore, observations reveal a general trend within the group {‘pullover’, ‘coat’, ‘shirt’}: an improvement in the accuracy of one class is accompanied by a corresponding decrease in another’s. The collective accuracy across these three classes appears to be relatively stable. The value hovers around $1/3$, implying that the model may be able to generally differentiate them from the dominant classes but fail to recognize the differences among them.
@@
Figure 5: The Class-wise Accuracy (%) During Training (FashionMNIST, Local Class Imbalance). X Axis: Rounds; Y Axis: Accuracy.
@@
While FashionMNIST is used as an example for ease of illustration, similar phenomena are observed in other datasets as well. For instance, in the MNIST dataset, samples from the weak classes ‘8’ and ‘9’ are more likely to be misclassified as {‘3’, ‘5’} and {‘4’, ‘7’}, respectively. In the CIFAR-10 dataset, the four weak classes {‘bird’, ‘cat’, ‘deer’, ‘dog’} are prone to confusion with one another. Interestingly, while these classes may not appear similar in their raw image forms, they exhibit similar high-level features that can make them challenging to distinguish for classification models.
III The Proposed Algorithm
III-A Overview

### v2

@@ Abstract.
Substantial efforts have been devoted to alleviating the impact of the long-tailed class distribution in federated learning. In this work, we observe an interesting phenomenon that certain weak classes consistently exist even for class-balanced learning. These weak classes, different from the minority classes in the previous works, are inherent to data and remain fairly consistent for various network structures, learning paradigms, and data partitioning methods. The inherent inter-class accuracy discrepancy can reach over 36.9% for federated learning on the FashionMNIST and CIFAR-10 datasets, even when the class distribution is balanced both globally and locally. In this study, we empirically analyze the potential reason for this phenomenon. Furthermore, a partial knowledge distillation (PKD) method is proposed to improve the model’s classification accuracy for weak classes. In this approach, knowledge transfer is initiated upon the occurrence of specific misclassifications within certain weak classes. Experimental results show that the accuracy of weak classes can be improved by 10.7%, reducing the inherent inter-class discrepancy effectively.
@@
Deep neural networks, federated learning, class imbalance, weak classes
1. Introduction
@@
Federated learning (FL) is a decentralized machine learning paradigm where multiple clients collaborate to train a global model without directly sharing their local data (McMahan et al., 2017). This approach preserves data privacy by retaining the data on local devices. However, a big challenge to FL is the varying performance across different classes due to the presence of unbalanced sample sizes (Zhang et al., 2023b). When the number of samples for different classes varies substantially, the global model often exhibits bias towards the majority classes, known as ‘head classes’. The minority classes, referred as ‘tail classes’, become underrepresented (Zhang et al., 2023a), thus leading to lower classification accuracy as depicted in Fig. 1(a). However, in this work, we show that ‘dominant’ and ‘weak’ classes consistently exist even when sample sizes are equalized and samples from each class are uniformly distributed across the clients (i.e., with a balanced class distribution both globally and locally). Unlike the minority classes discussed in prior studies, these weak classes are intrinsic to the data itself, rather than being dictated by sample size. The resulting inter-class accuracy discrepancy (ICD) is illustrated in Fig. 1(b).
@@
Fig. 2 presents the class-wise accuracy across several widely used benchmarks in the domain of FL. Each experiment is repeated five times, and the average result is presented. The ICD phenomenon identified in this study exhibits the following characteristics. 1) This phenomenon is universally present across different learning paradigms and is particularly pronounced in FL compared to centralized learning. 2) Neural networks trained on the same dataset tend to exhibit a very similar ICD, indicating that this phenomenon is irrelevant to the network structure. 3) This phenomenon persists no matter whether the local class distributions in FL is balanced or not. Notably, as local imbalance intensifies, particularly when samples from weak classes are concentrated within a limited number of clients, the inherent ICD becomes significantly exacerbated. Details will be elaborated upon in the subsequent sections.
@@
Figure 1. (a) The Conventional Long-tailed Problem; (b) Inherent Inter-class Discrepancies Observed under Balanced Class Distributions (both Globally and Locally).
@@
Figure 2. (a) The Class-wise Accuracy (with Mean Subtraction) Based on Different Learning Paradigms and Network Structures ($\Delta_{max}$ Refers to the Maximum Discrepancy after a Training Session; the Network Structures, such as VGG-9 (Simonyan and Zisserman, 2014; Wang et al., 2020) and ConvNets (McMahan et al., 2017), are Detailed in Section 4.1.) (b) Sample Distribution in Different Scenarios (Assuming a Total of 10 Clients).
@@
Although promising progress has been made in alleviating the inter-class discrepancy brought by class imbalance (Zhang et al., 2023a), these techniques cannot be directly applied to the inherent ICD problem. Specifically, current approaches usually focus on rebalancing class distributions by employing techniques such as over-sampling (Estabrooks et al., 2004) or data augmentation (Liu et al., 2020) for the minority classes. However, considering that the inherent ICD is not associated with the quantity of samples, these methods may not be sufficient to mitigate the problem effectively.
@@
Experiments indicate that a likely reason for this ICD phenomenon is the intrinsic similarities in high-level features between specific classes. This similarity poses a challenge for the classifier in accurately distinguishing data samples across these classes. In this work, we propose a federated partial knowledge distillation (PKD) method to improve classification accuracy for weak classes. Our method is inspired by the observation that an expert trained specifically on a group of confounding weak classes show greater proficiency in differentiating among them. This specialized model focuses solely on extracting and distinguishing the subtle feature differences among these classes and thus yields better accuracy than the model trained jointly on all classes. The term ‘partial’ in PKD is shown in several aspects: 1) An expert’s knowledge is confined to specific classes. It is tailored to identify a subset of data that shares analogous high-level features. 2) Knowledge distillation (KD) is triggered only when misclassifications between certain classes happens. Therefore, only a subset of the misclassified samples from these weak classes is considered in computing the KD loss during the training phase. 3) While there might be multiple experts, only up to one expert can be involved in the KD process for any given misclassified sample. Experimental results demonstrate that the proposed PKD method can effectively alleviate the inherent ICD issue by enhancing the classification accuracy of the worst performing classes. In summary, this work makes the following contributions:
@@
  • We highlight an interesting phenomenon that certain weak classes are always present, even in ideal scenarios where class distributions are balanced at both the global and local levels. This inherent variance in accuracy across different classes is especially pronounced in the context of FL, reaching up to 45.4% on the FashionMNIST dataset (Xiao et al., 2017) and 36.9% on the CIFAR-10 dataset (Krizhevsky, 2009), respectively.
@@
  • Experimental results show that the proposed method can effectively improve the classification accuracy of the weak classes by over 10.7% and 6.3% on FashionMNIST and CIFAR-10, respectively.
2. Inter-Class Accuracy Discrepancy
2.1. Discrepancy Induced by the Class Imbalance Problem
@@
ICD due to class imbalance. A primary reason for the variance in class-wise accuracy in deep learning is the imbalanced class distribution (He and Garcia, 2009; Buda et al., 2018). The unbalanced sample sizes lead to a dual-fold challenge: 1) Models tend to be biased towards the majority classes, leaving the minority classes under-represented. 2) The lack of data for minority classes makes it even harder to learn the characteristics of these classes (Zhang et al., 2023a). In federated learning, the potential causes for ICD become more complex. The presence of class imbalance at either the global level or within the local dataset of individual clients may diminish the model’s quality. Besides, the mismatch in class distributions between the cloud and the clients may also degrade the performance of the global model (Zhang et al., 2023b).
@@
Existing approaches. Existing solutions can be broadly classified into several categories. 1) Class re-balancing techniques: These methods focus on offsetting the negative impact of uneven class distribution in training samples. This can be achieved by either over-sampling the minority classes to ensure a more equitable presence in the training data (Estabrooks et al., 2004; Liu et al., 2022), or by re-weighting the training loss values to mitigate the influence of uneven positive gradients (Lin et al., 2020; Cui et al., 2019). 2) Data augmentation techniques: These methods seek to enlarge the sample size or enhance the sample quality of minority classes through data augmentation techniques (Chawla et al., 2002; Shuai et al., 2022; Kim et al., 2020). Compared to mere re-balancing methods, this approach can potentially enrich the diversity within the minority classes, thereby reducing the risk of overfitting. 3) Other methods include meta-learning (Ren et al., 2020), ensemble learning (Xiang et al., 2020; Cui et al., 2023), transfer learning (Xiang et al., 2020), and so on. Although significant progress has been achieved in reducing the ICD induced by class imbalance, such strategies may not be directly applicable to addressing the inherent ICD problem, as we will discuss below.
2.2. The Inherent ICD in This Work
@@
Differences from the class imbalance problem. While the weak classes show similar poor accuracy to the minority classes in the class imbalance problem, their underlying causes and potential solutions are distinct. First, the inherent ICD is not a result of uneven sample sizes across different classes. Weak classes consistently exist even in the ideal case where both global and local data are class balanced. Consequently, previous methods that rely on re-sampling local data (Shuai et al., 2022) or clients (Zhao et al., 2020) based on their sample sizes are not feasible in this context. Second, in contrast to minority classes, weak classes are not characterized by data scarcity or a lack of diversity. Experimental results suggest that the conventional data augmentation techniques (Devries and Taylor, 2017) fail to address the inherent ICD effectively, and in some cases, they might even intensify the inherent accuracy discrepancy.
@@
The inherent ICD phenomenon. Fig. 2(a) shows the class-wise accuracy achieved on three datasets. In each scenario, the total number of samples for each class is kept identical, ensuring that global class balance is maintained. We adopt several commonly used convolutional neural networks and train the models following different learning paradigms. Two local data partitioning strategies are adopted in FL for each dataset, as illustrated in Fig. 2(b). Our results indicate that this inherent ICD is a common phenomenon across various network structures, learning paradigms, and local data partitioning strategies. Most of the dominant classes and the weak classes, e.g., {‘0’, ‘1’} and {‘8’, ‘9’} in MNIST (Lecun et al., 1998), remain fairly consistent across various settings. The inter-class accuracy discrepancy is less marked for centralized learning and thus may go unnoticed. However, when it comes to FL, the discrepancy may increase significantly, e.g., reaching up to 45.4% for FashionMNIST. We also test the local class imbalance scenario using a pathological data partitioning strategy proposed in (McMahan et al., 2017). Specifically, each client is assigned samples from a single class to ensure a clear and straightforward comparison of the class-wise accuracy. As shown in Fig. 2(a), the same group of weak classes is observed, with the accuracy discrepancy further increasing to a significant 53.0% on the MNIST dataset. For more complex datasets like FashionMNIST and CIFAR-10, such discrepancy can reach up to over 73.9%. When evaluated under FL, CIFAR-10 exhibits a reduced discrepancy ($\Delta$) compared to FashionMNIST. This can be attributed to the degradation in overall model performance. For instance, under the extreme local class imbalance scenario, the average accuracy of VGG-9 (Simonyan and Zisserman, 2014; Wang et al., 2020) and ResNet-56 (He et al., 2016) notably drops below 20% on CIFAR-10.
@@
Fig. 3 shows the maximum and minimum class-wise accuracy at each round. While the overall accuracy rises smoothly over time, the difference between the maximum and minimum values continues to be evident. With a balanced class distribution globally and locally, the accuracy discrepancy on the FashionMNIST dataset is 28.9% when there are 10 clients. This value increases significantly as the number of clients grows to 100. In scenarios with local class imbalance, samples from weak classes are concentrated within a small subset of clients, thereby increasing the accuracy discrepancy to 91.2%. The minimum class-wise accuracy is close to zero, indicating that the final model might completely fail in specific classes.
@@
Figure 3. The Maximum, Average, and Minimum Class-wise Accuracy During Training.
Potential reasons for inherent ICD. Fig. 4 (c) maps the predictions to the original classes to which the data belongs (the models are trained once). It shows that the negative samples of a weak class are often assigned certain wrong labels. For the FashionMNIST dataset with a balanced local class distribution, 62% (=$\frac{85}{137}$) of the misclassified ‘T-shirt’ samples are identified as ‘shirts’ using the LeNet-5 model (Lecun et al., 1998). Conversely, around 42% (=$\frac{132}{316}$) of the misclassified ‘shirt’ samples are labeled as ‘T-shirts’. Another group of classes that can easily be confused with each other include {‘pullover’, ‘coat’, and ‘shirt’}. For instance, 23% of the misclassified ‘shirt’ samples are identified as ‘pullovers’, while 23% are mistaken for ‘coats’. That is to say, a total of 88% of the misclassified ‘shirt’ samples are incorrectly categorized as other weak classes. This phenomenon becomes particularly pronounced in scenarios with local class imbalance. Approximately 81% of the overall samples categorized under ‘pullover’ and ‘coat’ are misidentified as ‘shirt’ by the LeNet-5 model. A similar phenomenon is observed with other network models, as shown in Fig. 4 (c).
Figure 4. FashionMNIST: (a) Raw Samples and the Similarities in High-Level Features between each Pair of Classes; (b) Output Probabilities (Averaged over 1000 Samples per Class); (c) Prediction Results. For Simplicity, each Class is Assigned a Serial Number. (0: T-shirt, 1: Trouser, 2: Pullover, 3: Dress, 4: Coat, 5: Sandal, 6: Shirt, 7: Sneaker, 8: Bag, 9: Ankle Boot.)
Fig. 4(a) presents some misclassified samples from the two weak class groups, i.e., {‘T-shirt’, ‘shirt’} and {‘pullover’, ‘coat’, ‘shirt’}. It can be seen that these samples exhibit considerable similarities. Fig. 4(a) assesses the similarities in the high-level features ($f$) between each pair of classes. The weak classes within the two aforementioned groups exhibit a lower Euclidean distance compared to the others, indicating a higher degree of similarity. Therefore, it is challenging for the classifier to differentiate among these classes, leading to lower accuracy and reduced confidence in the model’s predictions. The probability distributions determined by the softmax function are depicted in Fig. 4(b). The figure displays the average probability distribution across all samples within each class. The t-SNE visualizations of feature vectors are presented in Fig. 5.
Figure 5. t-SNE Visualizations of Feature Vectors on (a) FashionMNIST and (b) CIFAR-10 (under Local Class-Balanced Scenario, 10 clients).
We further investigate how the classification accuracy for weak classes changes during training. As depicted in Fig. 6, the variance between two consecutive rounds can reach up to 80% under the local class-imbalanced scenario, even when the global accuracy has become relatively stable. Furthermore, observations reveal a general trend within the group {‘pullover’, ‘coat’, ‘shirt’}: an improvement in the accuracy of one class is accompanied by a corresponding decrease in another’s. The collective accuracy across these three classes appears to be relatively stable. The value hovers around $1/3$, implying that the model can generally differentiate weak classes from dominant classes but may fail to recognize the differences among the weak classes themselves.
Figure 6. The Class-wise Accuracy (%) During Training (FashionMNIST, under Local Class-Imbalanced Scenario). X Axis: Rounds; Y Axis: Accuracy.
@@
While FashionMNIST is used as an example for ease of illustration, similar phenomena are observed in other datasets as well. For instance, in the MNIST dataset, samples from the weak classes ‘8’ and ‘9’ are more likely to be misclassified as {‘3’, ‘5’} and {‘4’, ‘7’}, respectively. In the CIFAR-10 dataset, the four weak classes {‘bird’, ‘cat’, ‘deer’, ‘dog’} are prone to confusion with one another. Interestingly, while these classes may not appear similar in their raw image forms, they exhibit similar high-level features that can make them challenging to distinguish for classifiers.
3. The Proposed Algorithm
3.1. Overview